OpenStudy (anonymous):

Hooke's law. The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 14 cm when the weight is 2 kg, what is the distance when the weight is 5 kg?

OpenStudy (anonymous):

14/2=x/5 x=35

OpenStudy (anonymous):

so its 35 cm?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

love math is correct

OpenStudy (anonymous):

THX